OpenStudy (anonymous):

At the beginning of the week, Naveen has 87 trading cards. On Monday, he gets 9 more. On Tuesday, he gives 16 away. On Wednesday, he is given 3 by a friend. On Thursday, he sells all of his cards to a collector. The collector pays him a rate of c for every card that Naveen has. If Naveen earns $124.50 from the sale, what is c, the rate that the collector pays for each card?

OpenStudy (anonymous):

do anyone have the answer

OpenStudy (jebonna):

To work out the answer, work out how many cards he has when he goes to the collector, and divide the amount of money he gets by the amount of cards he has in the end of the days. :)

OpenStudy (bamboo888):

87+9=96-16=80+3=83. So he sells 83 cards. Divide 124.50 by 83=$1.50 each card.